Discussion
If P (A) and P (not A) are complementary events and P (A) = 0.15, then P (not A) =?
Answer: C
Complementary events cover every possibility between them, so their probabilities add to 1.
\[ P(\text{not } A) = 1 - P(A) = 1 - 0.15 = 0.85 \]
No comment is present. Be the first to comment.
Loading…
Post your comment
Please log in or register to post a comment.
Examination Categories
- SSC Exams
- Banking Exams
- Teaching Exams
- Civil Services Exams
- Railways Exams
- Engineering Recruitment Exams
- Defence Exams
- State Government Exams
- Police Exams
- Insurance Exams
- Nursing Exams
- Other Government Exams
- NRA CET
- PG Entrance Exams
- Campus Placement
- MBA Entrance Exams
- CUET and UG Entrance
- Judiciary Exams
- Regulatory Body Exams
- Software Development